OpenStudy (anonymous):

Hi, I need some help with a natural log problem: Determine the value of x if 6.03(1.23)^x=32.403. I have absolutely no clue how to re-format this. What would the first step be?

OpenStudy (tanner23456):

6.03(1.23)^x = 32.403 first divide the 6.03 1.23^x = 5.374 change this to a log log1.235.374=x Now just use the change of base formula to solve it. (log 5.374)/(log 1.23) = x

OpenStudy (anonymous):

Awesome, thanks!

OpenStudy (tanner23456):

no prob. glad to help